OpenStudy (anonymous):

The pentagon is composed of square ABCD and isosceles triangle BXC in which m

OpenStudy (kc_kennylau):

X is inside or outside ABCD?

OpenStudy (anonymous):

outside

OpenStudy (anonymous):

Is it like this? 

OpenStudy (anonymous):

yep thats what the picture looks like

OpenStudy (anonymous):

The area is the sum of the square and the triangle. Square: 4*4=16.

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

The area of the square?

OpenStudy (anonymous):

Its the area of the pentagon, would it be 24 then? or 20

OpenStudy (anonymous):

Notice that the shaded area is a fourth of that of the square and at the same time equal to that of the triangle. Thus the area of the triangle is 16/4 = 4 and the total area is 16+4 = 20. 

OpenStudy (anonymous):

oh I GOT IT. that makes sense thank you

OpenStudy (anonymous):

you're welcome